OpenStudy (anonymous):

if the first two elements of a number series are 1 and 2, and if each succeeding term is found by multiplying the two terms immediately preceding it, what is the fifth element of the series?

OpenStudy (anonymous):

\[a_1 = 1, a_2 = 2, a_3 = (1)(2) = 2, a_4 = (2)(2) = 4, a _5 = (4)(2) = 8\]